please help
-6x - 5y = 20
-y = 6x + 4

Answer by stanbon(48558)

1) -6x-5y=20
2) 6x+y -4
Add the two equations to get:
-4y=16
y=-4
Substitute into equation 2 to solve for "x":
6x+y=-4
6x-4=-4
x=0
Solution: x=0, y=-4
Cheers,
Stan H.